OpenStudy (anonymous):

Let a1 >sqroot 2, and let an+1=an/2+1/an for n=1,2,3.. a) Show that {an} is decreasing and an> sq 2 for n>or equal to 1 b) Why must {an} converge? c) find lim an as n approaches infinity

OpenStudy (anonymous):

is it \[a_{n+1}=\frac{a_n}{2}+\frac{1}{a_n}\]?

OpenStudy (anonymous):

yes and and a1 \[\sqrt{2}\]

OpenStudy (anonymous):

\[a_{n+1}=\frac{1}{2}a_n+\frac{1}{a_n},~a_1=\sqrt2\] To show it's decreasing, try to prove by induction: \[\begin{align*} n=1:&\\ &a_2=\frac{1}{2}a_1+\frac{1}{a_1}\\ &a_2=\frac{\sqrt2}{2}+\frac{1}{\sqrt2}\\ &a_2=\frac{2+\sqrt2}{2\sqrt2}<\sqrt2=a_1 \end{align*}\] So the claim that it's decreasing holds for the base case. Assume it holds for \(n=k\) and show that it holds for \(n=k+1\).

OpenStudy (anonymous):

And so it converges because the limit is less than 1?

OpenStudy (anonymous):

or equal to 1

OpenStudy (anonymous):

Well, no it's supposed to converge because it's monotonically decreasing (what you're showing in part (a)) and it's bounded (less than \(\sqrt2\), also a result of part (a)).

OpenStudy (anonymous):

If its bounded by less than \[\sqrt{2}\] then it will keep decreasing?

OpenStudy (anonymous):

The sequence is decreasing because you show that \(a_{n+1}<a_n\) for all \(n\) (as a result of the induction proof). You also know that \(a_n=\sqrt2\), so \(a_{n+1}<a_n=\sqrt2 \) for all \(n\). You know that every number in the sequence is positive, so the sequence is bounded between \(0\) and \(\sqrt 2\).

OpenStudy (anonymous):

(You still have yet to show that it's decreasing, by the way. I've started the proof for you, but it's only half done.)

OpenStudy (anonymous):

Ok that makes sense! Thanks so much!